### v1

@@ Appendix A MALS Galactic H i absorption line catalog
| 1 | Source_name | MALS name of the source (JHHMMSS.ss+DDMMSS.s) based on its right ascension and declination (J2000) in the continuum image from SPW_id specified in column 6. |
@@
| 2 | Pointing_id | The MALS pointing ID (JHHMMSS.ssDDMMSS.s) based on the position (J2000) of the central source in NVSS or SUMSS. |
@@
| 3 | Obs_date_U | The date and time (UTC) of the start of UHF-band observing block(s) in the format YYYY-MM-DDThh:mm. |
@@
| 4 | Obs_date_L | The date and time (UTC) of the start of L-band observing block(s) in the format YYYY-MM-DDThh:mm. |
| 5 | Obs_band | The observing band: L = L-band and U = UHF-band. |
@@
| 6 | SPW_id | This identifies the SPW corresponding to the spectrum. The possible values for extragalactic spectra in the barycentric frame are LSPW_i and USPW_i; here i goes from 0 to 14. For Galactic spectra in the LSRK frame, the possible values are LSPW_10G, LSPW_13G and LSPW_14G corresponding to H i 21-cm, OHSAT1 and OHMAIN line, respectively. |
| 7 | Maj_cbeam | The major axis (arcsec) of the restoring beam of the cube. |
| 8 | Min_cbeam | The minor axis (arcsec) of the restoring beam of the cube. |
| 9 | PA_cbeam | The position angle (degrees) of the restoring beam of the cube. |
@@
| 10 | Peak_pos | The right ascension and declination (J2000) of the pixel corresponding to flux density peak at which the spectrum is extracted. |
| 11 | Peak_pos_l | The Galactic longitude (degrees) of the Peak_pos. |
| 12 | Peak_pos_b | The Galactic latitude (degrees) of the Peak_pos. |
| 13 | Distance_pointing | The distance of the source (arcmin) from the pointing center. |
@@
| 14 | Scale_factor | The scale factor applied for the primary beam correction. The primary beam attenuation is corrected using the katbeam (version 0.1) model. |
| 15 | Peak_flux | The peak flux density (mJy beam-1) of the source. |
| 16 | Spec_rms | The spectral rms (mJy beam-1 channel-1) in the unsmoothed spectrum. |
@@
| 17 | Vis_flag | This is a boolean (True or False) indicating whether an absorption feature based on visual inspection is detected in the spectrum; Nan implies no visual inspection has been performed. |
@@
| 18 | n_clump | The number of well detached absorption features detected in the spectrum. |
@@
| 19 | n_comp | The total number of Gaussian components fit to the absorption features detected in the spectrum. |
| 20 | tau_max_val | The peak optical depth of the absorption feature. |
| 21 | tau_max_+err | The positive 1 error on tau_max_val. |
| 22 | tau_max_-err | The negative 1 error on tau_max_val. |
@@
| 23 | pos_max | The position of the tau_max_val; redshift (barycentric frame) for extragalactic absorption and in km s-1 (LSRK frame) relative to the rest frequency of H i 21-cm, OHSAT1 or OHMAIN line for Galactic absorption. |
| 24 | tau_int_val | The integrated optical depth (km s-1). |
| 25 | tau_int_+err | The positive 1 error on tau_int_val. |
| 26 | tau_int_-err | The negative 1 error on tau_int_val. |
| 27 | delta_v | The velocity width containing 50% of the tau_int_val. |
| 28 | comp_i_amp† | The amplitude of the Gaussian component fit to the absorption line. |
| 29 | comp_i_amp_+err† | The positive 1 error on Comp_i_amp. |
| 30 | comp_i_amp_-err† | The negative 1 error on Comp_i_amp. |
| 31 | comp_i_pos† | The position of the Gaussian component relative to pos_max (km s-1). |
| 32 | comp_i_pos_+err† | The positive 1 error on Comp_i_pos. |
| 33 | comp_i_pos_-err† | The negative 1 error on Comp_i_pos. |
| 34 | comp_i_sigma | The width ( in km s-1) of the Gaussian component. |
| 35 | comp_i_sigma_+err† | The positive 1 error on Comp_i_sigma. |
| 36 | comp_i_sigma_-err† | The negative 1 error on Comp_i_sigma. |

### v2

@@ Appendix A MALS Galactic H i absorption line catalog
1 | Source_name | MALS name of the source (JHHMMSS.ss+DDMMSS.s) based on its right ascension and declination (J2000) in the continuum image from SPW_id specified in column 6. |
@@
2 | Pointing_id | The MALS pointing ID (JHHMMSS.ssDDMMSS.s) based on the position (J2000) of the central source in NVSS or SUMSS. |
@@
3 | Obs_date_U | The date and time (UTC) of the start of UHF-band observing block(s) in the format YYYY-MM-DDThh:mm. |
@@
4 | Obs_date_L | The date and time (UTC) of the start of L-band observing block(s) in the format YYYY-MM-DDThh:mm. |
5 | Obs_band | The observing band: L = L-band and U = UHF-band. |
@@
6 | SPW_id | This identifies the SPW corresponding to the spectrum. The possible values for extragalactic spectra in the barycentric frame are LSPW_i and USPW_i; here i goes from 0 to 14. For Galactic spectra in the LSRK frame, the possible values are LSPW_10G, LSPW_13G and LSPW_14G corresponding to H i 21-cm, OHSAT1 and OHMAIN line, respectively. |
7 | Maj_cbeam | The major axis (arcsec) of the restoring beam of the cube. |
8 | Min_cbeam | The minor axis (arcsec) of the restoring beam of the cube. |
9 | PA_cbeam | The position angle (degrees) of the restoring beam of the cube. |
@@
10 | Peak_pos | The right ascension and declination (J2000) of the pixel corresponding to flux density peak at which the spectrum is extracted. |
11 | Peak_pos_l | The Galactic longitude (degrees) of the Peak_pos. |
12 | Peak_pos_b | The Galactic latitude (degrees) of the Peak_pos. |
13 | Distance_pointing | The distance of the source (arcmin) from the pointing center. |
@@
14 | Scale_factor | The scale factor applied for the primary beam correction. The primary beam attenuation is corrected using the katbeam (version 0.1) model. |
15 | Peak_flux | The peak flux density (mJy beam-1) of the source. |
16 | Spec_rms | The spectral rms (mJy beam-1 channel-1) in the unsmoothed spectrum. |
@@
17 | Vis_flag | This is a boolean (True or False) indicating whether an absorption feature based on visual inspection is detected in the spectrum; Nan implies no visual inspection has been performed. |
@@
18 | n_clump | The number of well detached absorption features detected in the spectrum. |
@@
19 | n_comp | The total number of Gaussian components fit to the absorption features detected in the spectrum. |
20 | tau_max_val | The peak optical depth of the absorption feature. |
21 | tau_max_+err | The positive 1 error on tau_max_val. |
22 | tau_max_-err | The negative 1 error on tau_max_val. |
@@
23 | pos_max | The position of the tau_max_val; redshift (barycentric frame) for extragalactic absorption and in km s-1 (LSRK frame) relative to the rest frequency of H i 21-cm, OHSAT1 or OHMAIN line for Galactic absorption. |
24 | tau_int_val | The integrated optical depth (km s-1). |
25 | tau_int_+err | The positive 1 error on tau_int_val. |
26 | tau_int_-err | The negative 1 error on tau_int_val. |
27 | delta_v | The velocity width containing 50% of the tau_int_val. |
28 | comp_i_amp† | The amplitude of the Gaussian component fit to the absorption line. |
29 | comp_i_amp_+err† | The positive 1 error on Comp_i_amp. |
30 | comp_i_amp_-err† | The negative 1 error on Comp_i_amp. |
31 | comp_i_pos† | The position of the Gaussian component relative to pos_max (km s-1). |
32 | comp_i_pos_+err† | The positive 1 error on Comp_i_pos. |
33 | comp_i_pos_-err† | The negative 1 error on Comp_i_pos. |
34 | comp_i_sigma | The width ( in km s-1) of the Gaussian component. |
35 | comp_i_sigma_+err† | The positive 1 error on Comp_i_sigma. |
36 | comp_i_sigma_-err† | The negative 1 error on Comp_i_sigma. |